Radisson Hotel had compiled a past year (2017) data on the numbers of rooms demanded by its clients as a result of room rate changes. During the off-peak season in May, a promotion was held and it discounted the room rate from $550 to $400. This promotion of room rate had brought in 700 room stays compared to 350 rooms stays before the room rate adjustment.                       

Calculate the price elasticity of demand of Radisson Hotel’s clients


1
Expert's answer
2022-01-24T14:19:05-0500

"E=\frac{\frac{4000}{550}}{\frac{700}{350}}=\frac{0.73}{2}=0.365" or 36.5%